Grade 8General Physics

Three capacitors each of capacitance 9 pF are connected in series. (a) What is the total capacitance of the combination? (b) What is the potential difference across each capacitor if the combination is connected to a 120 V supply?

prasanjeet kumar
12 Years agoGrade 8

1 Answer

Apoorva Arora
12 Years ago
\frac{1}{C_{eq}}=\frac{1}{C_{1}}+\frac{1}{C_{2}}+\frac{1}{C_{3}}
so C= 3 pF
for a parallel combination, the same voltage is distributed across each capacitor. So 120 V across all the three.
Thanks and Regards
Apoorva Arora
IIT Roorkee
askIITians Faculty